The force exerted by an electric charge at the origin on a charged particle at a point (x, y, z) with position vector r = x, y, z is F(r) = Kr/|r|3 where K is a constant. Find the work done as the particle moves along a straight line from (5, 0, 0) to (5, 3, 5).
Answer #1

representation of a line segment joining the points (a,b,c) & (l, min) is rlt) = 12-t).(a,b,c) +tuminy where te [0,1] therefoW = n kl 9t+25t) To (25+24+2/3/2 W it K(34t) at = Jo 195+ 3472)312 al (68t) dt W = 1 k So last 3442 3/2 put 25+94 +² = u & (6